### v1

# The Role of Human Creativity in the Presence of AI Creativity Tools at Work: A Case Study on AI-Driven Content Transformation in Journalism

Abstract.
As AI becomes more capable, it is unclear how human creativity will remain essential in jobs that incorporate AI. We conducted a 14-week study of a student newsroom using an AI tool to convert web articles into social media videos. Most treated the tool as a creative springboard, yet still had to edit many AI outputs. The tool enabled the team to publish successful content, receiving over 500,000 views. Yet creators sometimes treated AI as an unquestioned expert, accepting flawed suggestions. Editorial critique was essential to spot errors and guide creative solutions when AI failed. We discuss how AI’s inherent gaps ensure human creativity remains vital.
human creativity, generative AI, content transformation, creativity support tools, AI at work, journalism
ccs: Human-centered computing Empirical studies in HCI
1. Introduction
The integration of AI into creative processes is revolutionizing content creation, particularly in the fast-paced environment of newsrooms. As generative AI tools become increasingly prevalent, understanding the intricate interactions between these technologies and human creativity is crucial. What is the role of human creativity in the presence of AI creativity tools in the workplace? Does the role of human creativity change over time as people gain familiarity with AI tools? Despite the growing adoption of AI in creative domains, these questions remain underexplored. Although AI has some capacity for creation, creativity is a complex cognitive process involving many skills, including convergent thinking, divergent thinking, framing, analysis, critique, audience understanding, and iteration. AI may be able to assist with some or parts of these processes, but human creativity is likely to still play a large role in the full creative process.
To explore the role of human creativity in the presence of AI, this paper presents a 14-week study conducted within the social media team of a university journalism department’s newsroom, focusing on content creation for the TikTok platform. The team included five social media creators (all students), one editor (a faculty member), and two market researchers (research assistants), embodying a typical real-world newsroom structure. They used an AI tool designed to help creators transform web articles into engaging short video scripts and storyboards. Using the tool, the journalists read, selected, regenerated, and edited several possible AI-generated scripts before shooting and editing the videos themselves. Each journalist aimed to produce one new video a week and iterate on previous videos. The newsroom met weekly for the editor to review videos, request changes, and decide if a video was ready to publish. After 10 weeks, the market researchers conducted 42 interviews with members of the target audience to assess quality and gather suggestions for improvement. The study is complemented by comparative insights from a shorter case study involving two professional TV journalists.
We found that overall, creators used AI as a “creative springboard.” Four out of five creators consistently used the tool and they appreciated its ability to reframe news stories around characters—part of the novel social media format. Without the tool, it was difficult to balance information and entertainment. However, for final scripts, they often edited “about half” of the text for accuracy and clarity. Creators did choose to inject their own creative touches, especially in areas the tool did not support, such as adding visual elements like props or animations or using their acting ability to bring new emotions, attitudes, and depth to their characters. Creators sometimes chose not to use the tool when they felt confident about the story or had a specific creative inspiration. Overall, the team was successful in publishing 16 videos in 14 weeks, and amassing over 500,000 views.
Figure 1. An example of a news reel made during the newsroom study describes the three key facts of a story about a teacher shortage.
[…]
An example of a news reel made during the newsroom study describes the three key facts of a story about a teacher shortage. There are eight frames with two characters: a school board and a teachers union. thru dialog, they describe the problems the district faces with teacher shortages.
Human critical evaluation was essential in defending against AI errors. Surprisingly, creators often accepted poor AI suggestions, such as overly positive language and meaningless phrases. This was true of both student creators and experienced, professional TV journalists. Although journalists are experts in journalism, they are not experts in the new skill of making videos. Here, AI is perceived as an expert. Journalists defer to AI suggestions, which are coherent even if flawed in other ways. Additionally, AI can fail dramatically when its lack of context results in socially inappropriate suggestions. These flaws required creators’ critical intervention either to modify these narratives or to discard them entirely to uphold social appropriateness.
Creators’ relationship with the AI changed over the course of the 14-week study. Initially, they relied heavily on the AI’s outputs, but with time, creators demonstrated a growing tendency to integrate their creativity outside the AI’s direct outputs. When the team gathered feedback from target viewers, they found even more ways to make their content more engaging, beyond what the tool was offering. Gradual adaptation, editorial interventions, and learning from audience feedback over time all encouraged creators to continue to use their own creativity on top of the creativity support the tool offered.
2. Background and Related Work
Figure 2. ReelFramer is a human-AI co-creative system that supports journalists in creating news reels by transforming web articles into scripts and storyboards. ReelFramer enables the exploration of multiple narrative framings that span the infotainment spectrum and their foundational details. It also aids in visual framing for character and key visual detail design.
2.1. Studies of AI in a Work Context
Recent studies have mapped out the significant opportunities that generative AI brings to various industries, highlighting its potential to enhance productivity for a wide range of workspace tasks (Brynjolfsson et al., 2023; Noy and Zhang, 2023). It has shown promise in replacing both cognitive routine work (often considered drudgery) and cognitive non-routine work, including creative tasks (Bouschery et al., 2023). In an experimental study with a writing task carried out by 453 college-educated professionals, Noy and Zhang (2023) found that those using ChatGPT both saved time and increased quality. Two experiments with Boston Consulting Group (Dell’Acqua et al., 2023) show that consultants with access to ChatGPT were more productive and produced higher-quality ideas. Studies of AI coding tools show it leads to a 26% increase in weekly coding tasks completed over three companies (Cui et al., 2024; Peng et al., 2023).
However, various studies also express reservations about AI’s ability to fully replace human workers. A study of creativity support tools demonstrates that while the tools are helpful for specific tasks, they cannot solve problems entirely on their own (Tao et al., 2023)—especially for tasks involving core human skills, even in the long run (Long et al., 2024). AI support can even be problematic. Dell’Acqua et al. (2023) show that for difficult tasks, subjects using ChatGPT were about 20% less likely to make the correct recommendation, as many were led astray by ChatGPT’s convincing but incorrect analysis. Even when AI is correct, users face challenges incorporating AI results (Kim and Kang, 2024).
These limitations suggest that AI might not completely reshape entire industries but rather impact certain aspects of how people work and what they focus on (Randazzo et al., 2024). However, developing these new configurations requires learning which tasks AI can be trusted to complete and how best to incorporate its results.
2.2. AI for Creativity Support
The use of AI for creativity support, particularly to assist writing, has expanded rapidly in recent years, with applications across various domains, including journalism (Liu et al., 2022; Wang et al., 2024a; Petridis et al., 2023). AI writing assistants can provide scaffolding for writers by supporting different cognitive stages of the writing process (Flower, 1981), including framing (Hui and Sprouse, 2023; Wang et al., 2024a), idea generation (Clark et al., 2018; Petridis et al., 2023) to drafting (Gero et al., 2022; Zhang et al., 2023; Long et al., 2023) and revision (Lee et al., 2024; Gero et al., 2022; Wu et al., 2019). Lee et al. (Lee et al., 2024) provide a comprehensive design space for intelligent and interactive writing assistants that emphasizes the importance of considering task, user, technology, interaction, and ecosystem aspects when using AI to support writers. However, using AI in creativity also has risks; overreliance (Ashktorab et al., 2021; Kim et al., 2024; Vasconcelos et al., 2023; Zhang et al., 2024) on AI can cause people to accept poor AI answers, and AI ideation can actually reduce diversity across people (Anderson et al., 2024). Broadly, AI can be seen not as a solution, but as a design material that needs to be understood by human creators (Yildirim et al., 2022; Subramonyam et al., 2021).
A major creative application of AI lies in content retargeting and transformation—editing and repurposing text, image, video, or audio for new audiences, platforms, or goals. The key is to find the fundamental message to preserve while adapting the content into the new format. Previous works have explored how to support people with content transformation tasks in terms of video (Truong et al., 2021; Chi et al., 2021, 2022; Wang et al., 2024b, 2022). For example, AI systems can generate video podcast trailers by identifying salient scenes to summarize content while preserving its tone (Wang et al., 2024b), or shorten audio stories for social media consumption (Wang et al., 2022). However, creating effective short-form content poses unique challenges, particularly in planning and structuring narratives (Kim and Kim, 2024). Systems like ReelFramer (Wang et al., 2024a) address these hurdles by embedding genre-specific knowledge to guide writers in scripting videos that align with journalistic formats, exemplifying how AI can scaffold human creativity in multimedia journalism. By encoding domain expertise, such tools streamline the adaptation of long-form content into short, platform-optimized outputs.
2.3. News Reels
News reels help outlets reach younger viewers who favor TikTok and Instagram (Newman, 2022). Nearly half of major publishers now post on TikTok, and younger audiences are increasingly seeking news there (Newman et al., 2022). Although both news reels and traditional news aim to educate the audience on relevant information, the narrative forms and framings are very different. Traditional journalistic narratives are fact-driven and explain events in a serious tone. By contrast, news reels seek to balance information and entertainment by taking a more playful tone, in order to conform to the narrative structures common to social media videos. A common way of achieving this goal is to act out the story from the perspective of different characters that personify aspects of the story. For example, in a Washington Post reel [^1] https://www.tiktok.com/@washingtonpost/video/7147811605655473450, a character representing Hurricane Ian explains to another character representing the hurricane’s roommate why he is “visiting” Florida and all the trouble he might cause there. Even for a serious topic like hurricane destruction, news reels must strike a balance between informational content and entertainment value (Newman, 2022). Figure 1 contains an example of a news reel made in our newsroom study with a dialog between a School District character and a Teachers Union character acting out the news of a teacher shortage. As is typical of news reels, the creator is playing both roles.
Even within the role-play format, framing a news story for social media is challenging. In interviews conducted by the Reuters Institute in 2022 with news reel creators (Newman, 2022), many different approaches to creating reels have been found, but a key challenge is to balance between information and entertainment (Newman, 2022). If a serious topic is portrayed too light-heartedly, it may diminish the perceived seriousness of the issue (Davis et al., 2022). However, if a video does not grab a reader’s attention with something that is interesting and entertaining, they may scroll past it. Some news organizations take a conservative approach and stay closer to the TV journalism format, but others lean into the tropes of social media videos to blend news information into an engaging social media style. The challenge we address is to help journalists translate their articles into this style, skillfully blending information and entertainment to strike an appropriate tone.
2.4. ReelFramer AI Tool
ReelFramer (Wang et al., 2024a) is a creativity support system that uses generative text and image AI to scaffold the process of creating a script and storyboard for a news reel. Generative AI is helpful for exploring multiple narrative framings and their details, and generating scripts, character boards, and storyboards. The system is designed to be highly interactive, allowing people to make choices and guide the system at every step to ensure the news reels it produces are accurate, authentic, and achievable to execute.
The user starts by inputting the news article (see Figure 2). The system uses a large language model (LLM), GPT-4, to extract locations, people, and activities from the article. The user then selects one of the three narrative framings, expository dialogue, reenactment, or comedic analogy. Given a framing, the system uses the LLM to suggest foundational details like characters, plot, setting, and key information, which is also called premise in scriptwriting (Batty et al., 2017). The user can accept, regenerate, or edit these suggestions. Once the user finalizes a premise, the system automatically generates a script. Again, the user can accept, regenerate, or edit the script. Users are encouraged to explore many narrative framings, premises, and scripts. Once the user is happy with the script, the system provides visual framing to explore the characters’ visual design. It generates a character board that contains visual details of the costumes, props, and backgrounds to distinguish the characters using a text-to-image model, DALL-E 2. Based on the character board, it generates a storyboard that shows suggested emotions, actions, and dialogues for each character. Users can then make reels based on these materials.
3. Methods
3.1. Overview
We conducted three studies: 1) a 14-week study of how a small newsroom adopts a creativity support tool, 2) an audience reception study assessing the videos made with the tool, and 3) a 4-week study of how two professional TV journalists adopt a creativity support tool. We used iterative, exploratory methods (Creswell and Clark, 2017), guided by early insights from student journalists. Each study strand was analyzed independently yet informed by prior findings. Institutional review boards deemed the project exempt.
3.2. Organizational Cohort Field Study with Student Journalists
In the first study, researchers recruited advanced students from a university journalism program that runs a TV and print journalism media outlet called AnonymizedNews to be part of researchers’ simulated newsroom. As part of their existing program, the students report local area news, often publishing multiple stories a day with faculty guidance and approval. AnonymizedNews  has a fledgling TikTok channel, with fewer than 50 posts and no posts getting more than 3000 views. Simulated newsroom environments are common in journalism education spaces (Steel et al., 2007) as part of practical and experiential learning with both procedural (environment) and factual (task) authenticity (Chen et al., 2001). Choosing students in a university newsroom was fairly representative of professional newsrooms because they often hire recent graduates to run social media, due to their familiarity with social media and the demographic is aspired to reach.
Recruitment emails were sent to senior journalism classes offering an opportunity to learn about creating news reels with AI in a newsroom environment and to regularly publish videos under editorial guidance. Like an internship, this would be a paid opportunity. Six student journalists attended the first training session, and five participated in the full 14-week newsroom simulation. The size of this sample, along with attrition, is expected for longitudinal qualitative research (Saldana, 2003).
In this simulated newsroom environment, we had one researcher acting as editor, two researchers acting as market researchers, and five students given news stories each week to convert to reels, as if they were part of a digital news desk. The news stories were ones the newsroom had previously reported, sometimes by the student themselves. In weekly meetings, researchers acted as both participant observers  (Emerson et al., 2011), taking notes on how student journalists used the tool, and observant participants, interacting with the student journalists to offer the support an editor would provide (Seim, 2024), while allowing space for participants to share insights on production techniques with fellow student journalists as they would in a standard operating newsroom. Student journalists were regularly asked about their use or nonuse of the tool in the creation of their newsreels. The editor led group feedback discussions for every video, and gave specific feedback for iteration. After approval by the editor, newsreels were published to the AnonymizedNews  TikTok channel.
3.3. Audience Reception Study
To compare the quality of the student journalist-produced news reels to the other short form video news content (Shearer, 2024), we conducted qualitative interviews with typical short-form video audience members (Brennen, 2021). This allowed us to do two things: 1) access the relative quality of the student production, and 2) provide the student journalists with specific information to improve the audience reception of their videos. Questions covered how informative and entertaining the video was, their likelihood to watch the video if they came across it on their personal feed, and the likelihood of the video fitting in with content they typically see, from both qualitative and quantitative perspectives. Interviews took between 20 and 66 minutes, with an average of 35 minutes. 46 individuals signed up to participate, n=42 participants were interviewed.
Participants were shown six total short-form news videos. All six videos were systematically shown in a predetermined balanced design to evenly distribute order effects. Participants watched four videos created by the student journalists using ReelFramer on a range of topics. Two additional videos were added for comparison: one was professionally created by the Washington Post’s TikTok team to cover the trajectory of a hurricane, the other was created by a student in the AnonymizedNews  program but not in the study and not using the tool; it discussed a local basketball tournament.
Participants agreed to be interviewed in exchange for course credit as part of the university’s subject pool system. All participants were university students from across disciplines aged between 18-30, ($\mu$ = 20). Researchers recorded the interviews over Zoom and transcribed the videos before analysis began for axial themes (Saldana, 2003). The distribution of quantitative responses was plotted to compare the informativeness and fit of the news reels.
3.4. Qualitative Cohort Field Study with Practicing Journalists
To understand if there were differences in how professional journalists would adapt to the creativity support tool, researchers conducted a qualitative cohort field study with practicing journalists. In the summer of 2024, researchers recruited a cohort of journalists by posting in a social media group for multimedia journalists. Researchers informed participants we would hold four, one-hour collaborative sessions. 12 individuals completed a demographic questionnaire. Four individuals attended at least one training session, two produced at least one video using the tool. This sample size and these levels of attrition are not unexpected in longitudinal studies and do not diminish the benefit of the findings (Saldana, 2003). Similar to the student study, researchers acted in the capacity of newsroom editor offering critique and advice.
4. Study 1 Findings: Newsroom Field Study
4.1. AI Usage: AI as a Creative Springboard
Overall, the student group made 26 news reels, 16 of which were published by the editor. The 10 that were unpublished were still being iterated at the end of the session or had been abandoned (Section 4.2.2). Altogether, the 16 videos published by the team garnered over 500,000 views. Four of the five student journalists chose to use the tool for all or most of their videos. 13 of 16 published videos used ReelFramer in a significant manner to help achieve an output the editor deemed publishable.
4.1.1. Reasons for AI Tool Usage
Creators choose to use the AI tool a majority of the time—for 20 of 26 reels created. Creators who chose to use the tool often described it as “a creative springboard.” There were two main uses.
1. Selecting the narrative framing and premise. They all used the reframing features of ReelFramer to find the characters, premise, and key news facts in the web article that they could use to transform into a role-play script. They described this as a cognitively demanding task to “see the story” from a different narrative frame. ReelFramer’s suggestions were typically accurate, creating a reliable premise for a script.
2. Drafting multiple scripts. They used the scriptwriting writing features of ReelFramer to initiate a draft of the script. However, the users consistently reported editing “about half” of the AI-generated script. They found they had to improve many of the writing mechanisms, such as taking out awkward phrases, increasing clarity, or removing lines with too little content. Overall, it was a helpful starting point.
On the other hand, none of the students used the storyboarding features, as they did not find this part of the task cognitively demanding. They said they could “see” how the story would be filmed and apparently did not use the prop, costume, or gesture suggestions—often because they were constrained by their personal set of potential props, costumes, and acting ability.
4.1.2. Reasons and Risks of Not Using the Tool
Creators chose not to use the AI tool for 6 of the 26 reels made during the study. There were several legitimate reasons not to use the tool. However, choosing not to use the tool sometimes resulted in problems.
1. Strong understanding of the initial news story. Two times, a participant who had used ReelFramer previously chose not to use the tool when making a news reel for a story they had reported. Because they conducted the initial research and interviews and framed the original written story, it was easier to understand the key points in the story and imagine it from multiple perspectives, making it cognitively easier to reframe around a role-play narrative.
2. Task deemed easy. P5 made his first news reel with ReelFramer but chose to make the other three without it because he understood the framework well enough to do it on his own. However, this decision had trade-offs. When the editor reviewed the reels made without ReelFramer, she appreciated several creative additions that the tool would not have suggested (Section 4.2). However, the scripts were insufficient for publication—as they all lacked sufficient information about the story. The task of creating news reels is often more difficult than it looks; balancing between information and entertainment is hard, and the tool seems to help balance those competing design requirements.
3. Independent inspiration. P2 typically used ReelFramer, but for his fourth video, he had an inspiration for a role-play narrative and developed the framing and script on his own. However, he deviated from the role-play framing and wrote the script as a monologue, which the editor and other creators found confusing. Once he iterated to return it to role-play form, it was published. This incident also shows the risks of not using the scaffolding provided by the tool, although this creator was able to recover.
4. AI narratives inappropriate for a serious topic. P1 typically used ReelFramer, but for her second video, ReelFramer was not producing a narrative that was serious enough for her topic, which dealt with domestic abuse. ReelFramer’s “quirky” and “fun” narratives clashed with the gravity of the topic. This clash led her to create the news reel on her own. However, she had written the initial news story, so she had a strong understanding of it, and she took inspiration from the tool and followed a role-play format that delivered three news facts. The editor approved it for publication, and this news reel received by far the most views. There could be many reasons for this success—topic, timing, luck, or unknown quirks of the platform’s sharing algorithm. This is discussed in Section 5. Despite her success in making a news reel without the tool, she chose to return to using ReelFramer for all her future videos due to the cognitive support it provided.
Generally, the group was successful in their task, producing 16 news reels garnering over 500,000 views. The tool was treated as a creative springboard, assisting with cognitively demanding tasks of reframing a narrative, and writing a script that balanced entertainment and information. However, creators did not need a creative springboard when they had independent inspiration or when they had a strong understanding of the original web article. Additionally, when the tool failed to provide an appropriate framing or script, they had to fall back on their own creativity.
Figure 3. Creative additions made by humans that AI does and does not support.
4.2. Role of Human Creativity: Human Creativity Where AI Left off
To understand when and how creators chose to use their own creativity rather than relying on AI, we analyzed all their creative additions made to news reels created during the study that ReelFramer did not (or could not) make. Every week, when the editor reviewed news reels made by the creators (both published and unpublished), she noted all the instances when the creator did something creative outside of what ReelFramer suggested. After the study, we analyzed the types of tasks creators choose to innovate on, and found three clear groups of times when creators made creative additions.
4.2.1. Human Creativity to Improve AI Outputs
In the study, the editor noted a few tasks that were supported by AI, but for which the creators decided to improve with their own creativity by using their own framing or scriptwriting choices.
For framing, there were two times when a creator rejected AI character suggestions, and came up with characters that AI would not have: using an inanimate object (a gas pump, to talk about oil prices) or introducing characters not in the story (an Uber Driver, and a passenger to talk about a new bridge). For scriptwriting, one creator merged facts from two different web articles about the same topic (a local election) to tell a better story. ReelFramer only supported creating scripts from a single web article, so the creator had to add details to the script herself.
Creators frequently had to edit AI’s attempts at humor in scripts. AI often used puns, which creators did not like. In the middle of scripts, creators typically removed the attempted humor, but at the end of scripts, creators needed to come up with a new ending like saying something sarcastic. One creator added a curse word for emphasis said by a construction worker character. AI would never have suggested a curse word, and although the editor made him bleep it out before publishing, the audience still understood the effect, and it was a bold and creative choice. Although editing humor was common, the other creative framing and scripting edits were rare—occurring in only 4 of the 26 news reels.
4.2.2. Human Creativity for Major AI Failures
Major AI failures were rare—occurring in 3 of the 26 videos, but required significant creativity on the part of creators. In all three cases, AI made socially inappropriate suggestions because it lacked context of the story or the creators.
The first story was about new doula services offered for expecting mothers. The original web story featured quotes and reactions from women of color, since they are disproportionately at risk during pregnancies. AI suggested roleplaying characters from the story who were pregnant Black women. However, the creators did not feel like they could play this in a respectful way. But since the service is generally available for all people, they changed the characters to be people indirectly affected by the service (concerned family) and were able to report on the effects for women of all races. By creatively changing the characters, the creators were able to reframe the story in a socially appropriate way.
The second story was about domestic violence, which was a subject that was too serious for the lighthearted tone in the AI-suggested script. The creator chose to write the script herself, although she was able to creatively adapt the role-play framing by making the Domestic Women’s Shelter the expert explaining the dangers of emergency alert testing to the “naive” FCC—emergency alerts can expose hidden phones that domestic abuse victims often have.
The third, and most difficult case was when the team decided it was inappropriate to make a news reel out of a story. The article was titled “Women’s Club Basketball Discusses What Black History Month Means to Them”. None of the journalists identified as both Black and female, and thus role-play style narrative felt socially inappropriate. Additionally, they did not want to reframe the narrative around a different set of characters because it would displace the voices of the very people it was trying to center. Ultimately, they decided that not making the reel was the right decision.
Although major AI failures were relatively rare, they required human creativity to detect and address them.
4.2.3. Human Creativity Where AI Leaves off
In the study, the editor noted many tasks that were not supported by AI that creators decided to add their creativity to. This happened in 12 of 26 news reels (Figure 3). ReelFramer was designed to address the full creative process from framing to storyboarding. However, the creative process has no bounds, and the creators found many tasks to do creatively that the tool did not address.
There were four instances of creators adding personality to the characters, which was largely shown through acting ability. For example, making a teenager have an aggressively dismissive attitude about a new college program, to bring drama and realism to the character (as well as making it sound less like an advertisement).
There were three instances of using props in highly effective ways. First, using a real dog or baby instead of a doll, and second, in an episode about a soup kitchen, had a chef-like character actually making a dish and using a different cooking utensil in every scene to show progress. None of these were suggested by AI, but were things creators realized they had on hand.
The market research done in week 10 (Section 5) revealed that the videos would be more interesting if there were more movement, especially at the beginning. Thus, in the last week, creators added two more actions to videos – pouring maple syrup, jogging, and unwrapping food.
There were four other notable creative additions, like using animations of a truck for a story on monster trucks, adding sound effects, and adding colored lights and captions to reflect elements in a story.
Although the AI tool attempted to cover the entire creative process, creators found new dimensions where the tool left off, and they were able to add many of their own creative touches.
4.3. AI Risks: Need for Training Humans to Critique due to AI Overreliance
AI is widely known to produce errors—as are humans. The newsroom in this study had an editor who was in charge of quality control. This is typical of all newsrooms. Similar to typical editorial newsroom dynamics, creators submit drafts, receive critique and edits from the editor and have to iterate to produce a new draft. Iteration continues until the editor decides to publish the piece, or decide not to run it.
To understand the nature of the edits, we tracked two things: the magnitude of the error (major, minor or none), and the attribution of the error (whether the error was introduced by the AI or the human journalist). For error magnitude, major edits are those that required reshooting the entire news reel; minor edits required reshooting only part of the video (e.g., a localized scene or dialog edits); no edits meant it was ready for publication. Table 3 shows the magnitude of edits for each news reel’s first iteration.
We also tracked whether the errors were introduced by AI or by the creators, as judged by the editor from discussions with the creator. If the script contained incorrect information, and the script was drafted by AI and edited by a creator, the editor attributed this error to being introduced by AI. Although it should have been caught by a human, it might not have needed to be caught if it had not been introduced by AI. However, if the human writes a script with incorrect information, this error is attributed to the human.
Surprisingly, a vast majority of the errors were introduced by the AI tool. 11 of 15 major errors, and 15 of 19 minor errors were introduced by AI (Tables 1 and 2). These errors were accepted by the creators, and they filmed news reels with these errors in them. Although it is expected that AI will make errors, and that many people using AI on a deadline might simply accept AI content at face value, skimming the content and not checking it carefully. However, this case is different—creators actually had to speak the lines AI wrote, so they were fully cognizant of the dialog they were including. This was true for both student journalists and professional journalists (Section 6). It was surprising to the editor because being critical is an essential aspect of journalism training.
Major Issues (15) AI Attributed (11) Non-AI Attributed (4) too positive/promotional (4) slow paced (2) too long (2) confusing format - monologue (1) too few facts (2) biased narrative (1) unclear narrative (1) too wordy (1) abrupt intro (1)
Table 1. Major Issues: AI vs Non-AI Attributed issues in news reels created in the newsroom study.
Minor Issues (19) AI Attributed (15) Non-AI Attributed(4) awkward AI phrase (4) labels too long (1) confusing character names (3) added curse word (1) ending lacks polish (2) add relevancy to the present (1) too positive/promotional (1) missing facts (1) strays from news (1) tone shift (1) unclear/missing fact (2) unnecessary dialog (1)
Table 2. Minor Issues: AI vs Non-AI Attributed issues in news reels created in the newsroom study.
The most common major issue was clearly attributed to the AI tool—sounding too positive. The AI is trying to balance entertainment and information in a “quirky” way, and the stories should not sound sad and serious like traditional evening news anchors. However, the AI leaned too heavily toward a positive tone and this bias clashed with the content in different ways. For stories that were serious or sad (stories about homelessness, domestic violence, or climate change), the “quirky” tone was inappropriate for the topic. For many lighter stories—like a small town winning an award for being charming—AI’s positive tone was also a problem. The positive tone made the video feel like an advertisement rather than news. In all cases, the team had to find creative ways to fix the narratives to be more appropriate.
The most common minor issue clearly attributed to the AI tool (and accepted by creators) was including bizarre and awkward AI phrases. These stick out badly and are fairly localized changes to make, so the effort to change them should have been minimal. Phrases like “a little bird told me’’ are awkward AI phrases that were inappropriate for a news reel because all news should make some effort to attribute information to sources. Other awkward phrases include “a comedy of wonders” and distracting attempts at wordplay such as “a two-tiered system? is that kind getting rewarded with two kinds of deserts?” When asked why they included these phrases, the typical response was “I guess it seemed OK.”
This acceptance indicates an implicit trust in the AI, even when AI does a task in the creator’s realm of expertise. A crucial role of human creativity is to be critical of AI outputs. Even journalists who are trained to be critical may require additional training to learn to be critical in this new way.
4.4. Human Creative Role Over Time
We also tracked the magnitude and attribution of errors in the first iteration of all 26 news reels to track errors over time. As expected, every creator’s first video required major edits, but later videos needed more minor edits, or even no edits.
Participant Article 1 Article 2 Article 3 Article 4 Article 5 Article 6 P1 Major Minor Minor Minor No edits No edits AI errors 2 1 1 1 Non-AI errors 1 1 P2 Major Minor Major Major Major Minor AI errors 3 1 3 1 Non-AI errors 1 1 P3 Major Minor Minor Major / / AI errors 3 2 2 2 / / Non-AI errors / / P4 Major Minor Major Minor Minor No edits AI errors 1 1 2 1 1 Non-AI errors P5 Major Major Major Minor / / AI errors 2 / / Non-AI errors 1 1 1 1 / /
Table 3. Classification of the amount of editing needed for news reels made for each article and each participant. Edits were either major (requiring a full reshoot), minor (localized scene or line edits) or no edits. Errors are broken into those introduced by AI (and accepted by creators without realizing), or errors not introduced by AI. Cells filled with a slash indicate creators did not make news reels (some participants made more than others).
4.4.1. Learning the Limits of the Tool, and Reasserting Professional Standards
Early news reels were problematic because the news conveyed in them was unclear. This was likely due to a combination of problems: being too long, having meaningless dialog lines, and not containing enough facts. Additionally, when the tone was too positive, the reels felt like ads, not news, and when the roles of the characters were unclear, the video was hard to follow. All these problems could be attributed to AI, and creators’ early over-reliance on the tool. Through conversations with the editor and each other, they were able to overcome this. It seemed to be correlated with both more group and editorial critique, as well as students having a better mental model of the strengths and weaknesses of the tool through repeated usage. The editor pointed out that “at some point [the students] were able to anticipate what the tool was going to suggest (to a point) because they understood how it was translating between genres [web stories to reels].”
In the early videos, students often “miss the heart of the story.” For example, in a video about falling gas prices, the student “missed the main connection between the winter months and the change in gas prices”. Many exchanges in the early videos are wordy and contentless and do not help convey the news. It was important for students to understand the “news value” so the videos focused on the information rather than entertainment. The editor said that the students’ “learning/improvement happened when they made connections between the tool and their occupational standards.” Balancing between news and entertainment was one of the goals of the tool, and whereas it did help, it was still important for the creators to have their own understanding of this, so they could make good edits to what the tool was suggesting.
4.4.2. Learning from market research
In week 10 of the newsroom study, the team was given a report from the market research team detailing ways their news reels could be improved. This included more attention-grabbing elements early in the video, and more character motion-like actions within the scene (Section 5). These suggestions provided valuable insights because these were all emerging trends appearing in non-news-based social media videos that had not yet become integrated with news reels and were imbued into ReelFramer. Since news reels have to compete for attention with all the other reels, they will eventually have to adopt these trends to maintain their audience.
Working as a team, the creators brainstormed relevant actions they could put in their current videos and found several compelling ways to do this. For an article about tapping maple syrup trees early due to global warming, the video started with the character pouring syrup on waffles while talking. Another video about the opening of baseball season had several characters using props and doing simple, yet interesting actions like unwrapping food. A third added fun animation of a truck for a story about a monster truck rally.
The market research feedback was insightful, relevant and easy to apply. It was also outside the scope of AI, which cannot keep up with emerging trends.
5. Study 2 Findings: Audience Reception
Figure 4. Violin plots showing the distribution of ratings of videos as entertaining on a 1-7 scale. P is the professionally-created video, S the student-created video without using the tool and RF1-4, four videos created using the tool. The small letters labelling the bars identify groups of videos that are not statistically significantly different according to Tukey’s HSD. Multiple letters indicate that video is in two groups. Specifically, there is not a significant difference in the rating between videos P, S, RF2 and RF4, nor between video S and RF1-4 but the differences between video P and RF1 and RF3 are significant.
Figure 5. Violin plots showing the distribution of ratings of videos as entertaining. Labels are the same as Figure 4.
The main challenge in a news reel is to balance information and entertainment. We test how informative and entertaining news reels created in the study (using ReelFramer) were, compared to a popular professional news reel, and a news reel produced by a student at the same university but who was not in the study and was not exposed to ReelFramer.
First, we assessed how entertaining the news reels made during the story with ReelFramer were compared to the two baseline reels. According to a one-way ANOVA test, there was a significant difference in how entertaining the videos were ($F(5,229)=3.3$, $p=0.007$). Half of the videos produced using the tool were rated as as entertaining as the professionally made video. The non-study student’s video was also as entertaining as the popular video (Figure 4). However, it was rated less information than 3 of 4 ReelFramer videos.
Next, we assess how informative the news reels produced in the study are to the popular video and the non-study student reel. A one-way ANOVA test (not assuming equal variance in the groups) shows participant ratings of the video’s informativeness differed significantly ($F(5,108.18)=7.7$, $p<0.001$). Three-quarters of news reels produced in the study using ReelFramer were rated as “as informative as” the professionally produced video (Figure 5).
These results indicate that ReelFramer reels successfully balanced information and entertainment much of the time. They were competitive with the popular professional news reel, and it was better than the non-study student reel because they were more informational.
To assess audience acceptance of the news reels, we asked subjects whether they would watch each of the videos. Roughly three-quarters of subjects said they would watch the professionally produced video, 56% would watch the videos created with the tool and 45% would watch the student video produced without the tool. This difference was statistically significant ($\chi^{2}(5)=14.6$, $p=0.01$). Roughly half said they would share the professionally produced video, 40% would share the study videos, and only a quarter would the non-study student video. This difference was statistically significant ($\chi^{2}(5)=11.6$, $p=0.04$) There was no statistically significant difference in the perceived “fit” of videos on participant feeds ($\chi^{2}(5)=2.52$, $p=0.77$), though the student-produced video had a lower fit than other videos. Qualitative data suggests that this has to do with the subjects they prefer to watch (pets, sports, etc.) rather than quality. These results suggest that the videos produced with ReelFramer were closer to the professionally produced video in terms of acceptability than the non-study student video.
After conducting audience reception interviews, researchers identified strategic and manageable production changes they suggested to student journalists. Audience members indicated they would like more attention-grabbing elements early in the video and adding moving character movement within the scene. Student journalists wove this feedback into their subsequent productions and asked for more from the tool to match audience expectations. Similar to the way a newsroom would operate, journalists had to option of bringing market research into their production, continuing to iterate on their creative process.
6. Study 3 Findings: Professional Journalists
To supplement the results of the newsroom study done with student journalists, we conducted a 4-week study with two professional TV journalists from local news stations in the US.
6.1. Reasons for AI Tool Usage
Both journalists made 3 news reels each, for a total of 6 newsreels. They both chose to use the AI tool for all their videos. Similar to the newsroom study, they used the AI tool for framing and script writing, but not for storyboarding—they described framing and script writing as feeling “very different” than their training, and difficult to “get started with.” They also said that they had to edit “about half” the AI scripts for clarity, correctness, and smooth language.
6.2. Role of Professional Human Creativity: Human Creativity Where AI Leaves off
In four of the six videos, the journalists made notable creative additions. This included some of the same categories seen in the student study—adding personality to characters through acting and adding visual effects. However, they also introduced new elements. Two videos included real footage as background imagery. One video used an audio track. One video took the role-play premise to an extreme and introduced a time-lapse—where a citizen called city services over the course of 10 years to get a fallen tree removed (the city representative promised to “get back to them” in a robotic voice in every call). With professional journalism training and resources, it is possible to add even more creativity where the AI tool leaves off.
Of the six videos in the study, none experienced the major AI failures that forced students to make the news without the tool. However, AI did create some problems. Participant Pro 1 made a video discussing a new scam. The character introducing the scam was the Better Business Bureau (BBB). Although the typical TV news audiences know what the BBB is, Gen-Z and even Millennial audiences are less familiar with and are confused about the connection between the BBB and scams. AI was not aware of this gap in audience knowledge, but the editor was. But that decision still required much discussion and consideration.
6.3. AI Risks: Even Professionals Need to Hone Critique of AI
Major Issues (6) AI Attributed (4) Non-AI Attributed(2) unnecessary dialog (2) slow paced (1) too few facts (1) added timeline narrative unclear (1) too long (1)
Table 4. Major Issues: AI vs Non-AI Attributed issues in news reels created in the professional study.
Minor Issues (8) AI Attributed (5) Non-AI Attributed(3) confusing character names (3) distracting music (2) awkward AI phrase (2) should not name specific unrelated brands (1)
Table 5. Minor Issues: AI vs Non-AI Attributed issues in news reels created in the professional study.
Despite years of journalistic practice, the professionals still made many errors. Four of the six major errors and five of eight minor errors were introduced by AI and not caught by professionals before filming (see Tables 4 and 5). Generally, they were the same types of errors seen in the student study. This included major errors like having unnecessary and contentless dialog (“I can’t thank you enough”, “We’re so happy to help”), having too few facts, and being too long. As well as minor issues, like using awkward AI phrases (“How about I come and start training my own railings?”). The editor thought these should have been glaring errors to catch. Upon viewing the iteration of the video containing that dialog, the editor asked if “training railings” was a concept the professional journalist knew (perhaps a local phrase). The journalist said no, “The AI said it; I thought it was a thing.” Despite years of journalism training—a field that prides itself on being critical—these professionals were imperfect AI critics.
6.4. Professional Creative Role Over Time
Similar to the student study, news reels made by professional journalists still required major edits. However, both professionals required no edits by their third video (see Table 6). In both videos needing major edits, the errors were similar to the students’—the news content was unclear due to a variety of errors, like contentless dialog, and unclear roles. They did have to be coached by the editor to correct AI, but once the editor pointed that out, they adapted almost immediately. This is likely because they had a wealth of journalism training and understanding of “news value”, and thus faster to integrate their news training into their use of the tool.
Participant Article 1 Article 2 Article 3 Pro1 Major Minor No edits AI errors 1 3 Non-AI errors 3 1 Pro2 Minor Major No edits AI errors 1 3 Non-AI errors 1
Table 6. Classification of the amount of editing needed for news reels made by professional journalists.
7. Discussion and Conclusion
7.1. Experts Value AI Creative Support and Use Their Creativity Where AI Leaves Off
We found that when AI presents a solution, that solution, unless obviously flawed, is generally adopted. The user then applies their creativity to the parts of the project not solved by AI. There may be something to be learned from other areas of design. Lars Lerup, in a book called Building the Unfinished, observed that the physical environment will change over time: for example, we renovate buildings (Lerup, 1977). Instead of designing something that is a finished product, it might be better to build it in a way that encourages or facilitates later modification. A contemporaneous and related idea from architecture has to do with planning and building incrementally rather than all at once (Alexander, 1977).
AI leaves many things unfinished. In the newsroom study, AI failed in spectacular ways, leaving people to complete the task. AI also does not know what the next trend will be. In the newsroom study, we saw how the results from the market research were immediately used by journalists to keep up with social media engagement trends that AI was unaware of. Creativity is about discovering what comes next, and so AI has not shown abilities to do that.
With all automated tools there is a fear that human skills will atrophy from disuse. However, in this study, we saw rampant use of creativity, especially where the tool left off. This incompleteness of AI could be thought of as a feature rather than a bug—it is an opportunity to keep us engaged with the tool. Future work might explore the different ways of leaving an artifact unfinished and their effects on the quality of the end product, the overall level of productivity, and the long-term effect on human skill.
7.2. Creating Vigilance to AI Errors
AI makes errors. Often it makes obvious errors like using totally made up phrases. It was surprising to see that journalists, even professionals, filmed videos without fixing them. There seems to be a pervasive assumption that AI is an expert, even when AI is operating in a field in which the human is the expert and the AI is meant to play a supporting role.
Whatever the reason for their deference, people need to develop a vigilance towards AI errors. AI is so easy to generate that it seems to lull us into System I reasoning (Kahneman, 2011), where we think fast and mostly go with the flow. However, when working with AI, we need to develop ways to help people shift into System II reasoning, where we think slowly and think critically. System I is our natural state because it will take more than just willpower to develop this vigilance.
Teams and organizations will have to design ways to avoid falling for AI errors. Editors are needed to be critical; editors will need to learn the new flaws that AI is introducing into creative work. For example, AI is known to lean toward cliches in creative domains  (Chakrabarty et al., 2024; Mirowski et al., 2024), known to make up phrases (as shown here), known to lack social context, and known to be bad at simple math (Schick et al., 2023). When possible, it might also be helpful to constantly track what text or images were AI-generated so editors know where to focus their attention.
7.3. Limitations and Future Work
These studies attempt to explore AI at work in a newsroom, but we only studied one news team—and it was one that we assembled for this purpose. Although it had many of the traits of a real small-town newsroom, there are limitations in how much we can generalize to all newsrooms. Although the study lasted 14 weeks, there is no telling what would have happened if the study had continued after those 14 weeks. While we did not see any skill atrophy in the 14-week period, perhaps we would have after a year.
This paper focused on one particular tool for one particular social media phenomenon, but different tools and trends may have different benefits and challenges. However, factors such as accepting AI errors seem to be pervasive problems. ReelFramer as a tool is certainly not perfect. If its workflow had been better, perhaps creators would not have had so many opportunities to be creative when it failed. However, the most important failures were the ones when the new reel form did not fit an article—which was not a property of the tool, but of the narrative form.
It would be interesting future work to see if people were still using the tool in a year or whether they had embodied the news reel form so well the tool no longer brought them cognitive benefit. Or perhaps they might even find it too limiting over time, as they started to understand the design space and perhaps start to innovate on it. It would also be interesting to study how the team might improve the tool for themselves. Since the tool is basically a structured prompting technique, the creators could easily edit the tool by simply editing the prompts in the text. Another form of creativity would be to use prompts within the tool to innovate on the form.

### v2

@@
 We conducted a 14-week study of a student newsroom using an AI tool to convert web articles into social media videos.
-Most treated the tool as a creative springboard, yet still had to edit many AI outputs.
-The tool enabled the team to publish successful content, receiving over 500,000 views.
-Yet creators sometimes treated AI as an unquestioned expert, accepting flawed suggestions.
-Editorial critique was essential to spot errors and guide creative solutions when AI failed.
-We discuss how AI’s inherent gaps ensure human creativity remains vital.
+Most creators treated the tool as a creative springboard, not as a completion mechanism.
+They edited the AI outputs.
+The tool enabled the team to publish successful content that received over 500,000 views.
+Human creativity remained essential: after AI produced templated outputs, creators took ownership of the task, injecting their own creativity, especially when AI failed to create appropriate content.
+AI was initially seen as an authority, due to creators’ lack of experience, but they ultimately learned to assert their own authority.
 human creativity, generative AI, content transformation, creativity support tools, AI at work, journalism
-ccs: Human-centered computing Empirical studies in HCI
 1. Introduction
@@
 Does the role of human creativity change over time as people gain familiarity with AI tools?
-Despite the growing adoption of AI in creative domains, these questions remain underexplored.
-Although AI has some capacity for creation, creativity is a complex cognitive process involving many skills, including convergent thinking, divergent thinking, framing, analysis, critique, audience understanding, and iteration.
-AI may be able to assist with some or parts of these processes, but human creativity is likely to still play a large role in the full creative process.
+Although AI has some capacity for creation, creativity is a complex cognitive process involving many skills, including divergent and convergent thinking, framing, analysis, critique, audience understanding, and iteration.
+AI may be able to assist with some or parts of these processes, but human creativity is likely to still play a role in the creative process.
 To explore the role of human creativity in the presence of AI, this paper presents a 14-week study conducted within the social media team of a university journalism department’s newsroom, focusing on content creation for the TikTok platform.
@@
 The newsroom met weekly for the editor to review videos, request changes, and decide if a video was ready to publish.
-After 10 weeks, the market researchers conducted 42 interviews with members of the target audience to assess quality and gather suggestions for improvement.
-The study is complemented by comparative insights from a shorter case study involving two professional TV journalists.
-We found that overall, creators used AI as a “creative springboard.” Four out of five creators consistently used the tool and they appreciated its ability to reframe news stories around characters—part of the novel social media format.
-Without the tool, it was difficult to balance information and entertainment.
-However, for final scripts, they often edited “about half” of the text for accuracy and clarity.
-Creators did choose to inject their own creative touches, especially in areas the tool did not support, such as adding visual elements like props or animations or using their acting ability to bring new emotions, attitudes, and depth to their characters.
-Creators sometimes chose not to use the tool when they felt confident about the story or had a specific creative inspiration.
-Overall, the team was successful in publishing 16 videos in 14 weeks, and amassing over 500,000 views.
-Figure 1. An example of a news reel made during the newsroom study describes the three key facts of a story about a teacher shortage.
-[…]
-An example of a news reel made during the newsroom study describes the three key facts of a story about a teacher shortage.
-There are eight frames with two characters: a school board and a teachers union.
-thru dialog, they describe the problems the district faces with teacher shortages.
-Human critical evaluation was essential in defending against AI errors.
-Surprisingly, creators often accepted poor AI suggestions, such as overly positive language and meaningless phrases.
-This was true of both student creators and experienced, professional TV journalists.
-Although journalists are experts in journalism, they are not experts in the new skill of making videos.
-Here, AI is perceived as an expert.
-Journalists defer to AI suggestions, which are coherent even if flawed in other ways.
-Additionally, AI can fail dramatically when its lack of context results in socially inappropriate suggestions.
-These flaws required creators’ critical intervention either to modify these narratives or to discard them entirely to uphold social appropriateness.
-Creators’ relationship with the AI changed over the course of the 14-week study.
-Initially, they relied heavily on the AI’s outputs, but with time, creators demonstrated a growing tendency to integrate their creativity outside the AI’s direct outputs.
-When the team gathered feedback from target viewers, they found even more ways to make their content more engaging, beyond what the tool was offering.
-Gradual adaptation, editorial interventions, and learning from audience feedback over time all encouraged creators to continue to use their own creativity on top of the creativity support the tool offered.
+After 10 weeks, the market researchers conducted 42 interviews with members of the target audience to gather suggestions for improvement.
+Figure 1. A news reel made during the student study covering local news about a teacher shortage.
+Overall, the team was successful in publishing 16 videos in 14 weeks and amassing over 500,000 views.
+AI was far from perfect—creators had to edit “about half” of the AI scripts for accuracy and clarity.
+However, AI was helpful both as scaffolding when learning the task, and as a creative springboard (Shneiderman, 2007) even after they understood the task.
+Additionally, the tool was not strictly necessary—creators opted not to use the tool when it was not working well for a story or if they had a specific creative inspiration.
+In addition to correcting AI, creators also intervened to inject their own creativity in many parts of the process.
+Whereas AI could only produce outputs, the creators took ownership over the task, using their judgment and creativity to maximize its value to audiences, rather than just adapting a story to a template as AI did.
+The creators’ relationship with AI changed over the course of the 14-week study.
+Initially, they relied heavily on the AI’s outputs—even accepting its blatant errors without question.
+However, through critical reflection and iteration, creators were able to reassert themselves.
+AI was initially seen as “an authority” on news reels, due to creators’ lack of experience and AI’s mystique of intelligence.
+But ultimately, users were better able to balance AI’s authority and their own.
 2. Background and Related Work
 Figure 2.
-ReelFramer is a human-AI co-creative system that supports journalists in creating news reels by transforming web articles into scripts and storyboards.
-ReelFramer enables the exploration of multiple narrative framings that span the infotainment spectrum and their foundational details.
-It also aids in visual framing for character and key visual detail design.
+ReelFramer is a human-AI co-creative system that supports journalists in creating news reels from print articles.
+Users input the text of a news article, and accept or reject AI suggestions for news information, narrative framing elements, script, visual framing, and a storyboard.
 2.1. Studies of AI in a Work Context
 Recent studies have mapped out the significant opportunities that generative AI brings to various industries, highlighting its potential to enhance productivity for a wide range of workspace tasks (Brynjolfsson et al., 2023; Noy and Zhang, 2023).
-It has shown promise in replacing both cognitive routine work (often considered drudgery) and cognitive non-routine work, including creative tasks (Bouschery et al., 2023).
-In an experimental study with a writing task carried out by 453 college-educated professionals, Noy and Zhang (2023) found that those using ChatGPT both saved time and increased quality.
-Two experiments with Boston Consulting Group (Dell’Acqua et al., 2023) show that consultants with access to ChatGPT were more productive and produced higher-quality ideas.
-Studies of AI coding tools show it leads to a 26% increase in weekly coding tasks completed over three companies (Cui et al., 2024; Peng et al., 2023).
+It has shown promise in replacing both cognitive routine work (often considered drudgery) and cognitive non-routine work, including creative tasks (Bouschery et al., 2023), professional writing tasks (Noy and Zhang, 2023), professional software development (Cui et al., 2024; Peng et al., 2023), and professional analysis work (Dell’Acqua et al., 2023).
 However, various studies also express reservations about AI’s ability to fully replace human workers.
@@
 AI support can even be problematic.
-Dell’Acqua et al.
-(2023) show that for difficult tasks, subjects using ChatGPT were about 20% less likely to make the correct recommendation, as many were led astray by ChatGPT’s convincing but incorrect analysis.
+(Dell’Acqua et al., 2023) shows that for difficult tasks, subjects using ChatGPT were about 20% less likely to make the correct recommendation, as many were led astray by its convincing but incorrect analysis.
 Even when AI is correct, users face challenges incorporating AI results (Kim and Kang, 2024).
 These limitations suggest that AI might not completely reshape entire industries but rather impact certain aspects of how people work and what they focus on (Randazzo et al., 2024).
-However, developing these new configurations requires learning which tasks AI can be trusted to complete and how best to incorporate its results.
 2.2. AI for Creativity Support
 The use of AI for creativity support, particularly to assist writing, has expanded rapidly in recent years, with applications across various domains, including journalism (Liu et al., 2022; Wang et al., 2024a; Petridis et al., 2023).
-AI writing assistants can provide scaffolding for writers by supporting different cognitive stages of the writing process (Flower, 1981), including framing (Hui and Sprouse, 2023; Wang et al., 2024a), idea generation (Clark et al., 2018; Petridis et al., 2023) to drafting (Gero et al., 2022; Zhang et al., 2023; Long et al., 2023) and revision (Lee et al., 2024; Gero et al., 2022; Wu et al., 2019).
-Lee et al.
-(Lee et al., 2024) provide a comprehensive design space for intelligent and interactive writing assistants that emphasizes the importance of considering task, user, technology, interaction, and ecosystem aspects when using AI to support writers.
+AI writing assistants can provide scaffolding for writers by supporting different cognitive stages of the writing process (Flower, 1981), including framing (Hui and Sprouse, 2023; Wang et al., 2024a), idea generation (Clark et al., 2018; Petridis et al., 2023) to drafting (Gero et al., 2022; Zhang et al., 2023; Long et al., 2023; Menon et al., 2024) and revision (Gero et al., 2022; Wu et al., 2019).
+(Lee et al., 2024) provides a comprehensive design space for intelligent and interactive writing assistants that emphasizes the importance of considering task, user, technology, interaction, and ecosystem aspects when using AI to support writers.
 However, using AI in creativity also has risks; overreliance (Ashktorab et al., 2021; Kim et al., 2024; Vasconcelos et al., 2023; Zhang et al., 2024) on AI can cause people to accept poor AI answers, and AI ideation can actually reduce diversity across people (Anderson et al., 2024).
@@
 The key is to find the fundamental message to preserve while adapting the content into the new format.
-Previous works have explored how to support people with content transformation tasks in terms of video (Truong et al., 2021; Chi et al., 2021, 2022; Wang et al., 2024b, 2022).
-For example, AI systems can generate video podcast trailers by identifying salient scenes to summarize content while preserving its tone (Wang et al., 2024b), or shorten audio stories for social media consumption (Wang et al., 2022).
-However, creating effective short-form content poses unique challenges, particularly in planning and structuring narratives (Kim and Kim, 2024).
-Systems like ReelFramer (Wang et al., 2024a) address these hurdles by embedding genre-specific knowledge to guide writers in scripting videos that align with journalistic formats, exemplifying how AI can scaffold human creativity in multimedia journalism.
-By encoding domain expertise, such tools streamline the adaptation of long-form content into short, platform-optimized outputs.
+Previous works have explored how to support people with content transformation tasks in long-form and short-form video creation (Truong et al., 2021; Chi et al., 2021, 2022; Wang et al., 2024b, 2022).
+In particular, short-form content requires restructuring narratives to fit new constraints and audience expectations (Kim and Kim, 2024).
 2.3. News Reels
-News reels help outlets reach younger viewers who favor TikTok and Instagram (Newman, 2022).
-Nearly half of major publishers now post on TikTok, and younger audiences are increasingly seeking news there (Newman et al., 2022).
-Although both news reels and traditional news aim to educate the audience on relevant information, the narrative forms and framings are very different.
-Traditional journalistic narratives are fact-driven and explain events in a serious tone.
-By contrast, news reels seek to balance information and entertainment by taking a more playful tone, in order to conform to the narrative structures common to social media videos.
-A common way of achieving this goal is to act out the story from the perspective of different characters that personify aspects of the story.
-For example, in a Washington Post reel [^1] https://www.tiktok.com/@washingtonpost/video/7147811605655473450, a character representing Hurricane Ian explains to another character representing the hurricane’s roommate why he is “visiting” Florida and all the trouble he might cause there.
-Even for a serious topic like hurricane destruction, news reels must strike a balance between informational content and entertainment value (Newman, 2022).
-Figure 1 contains an example of a news reel made in our newsroom study with a dialog between a School District character and a Teachers Union character acting out the news of a teacher shortage.
+News reels help outlets reach younger viewers who favor TikTok and Instagram (Newman, 2022; Newman et al., 2022).
+Whereas traditional journalistic narratives are serious and fact-driven, news reels often use a playful tone that balances between information and entertainment, which meets expectations for social media videos (Newman, 2022).
+A common way to reframe news for reels is to create a role-play.
+An actor/journalist creates a dialogue between two characters related to the news story and acts out the perspective of each character.
+For example, in a Washington Post reel, a character representing Hurricane Ian explains to a roommate character why he is “visiting” Florida and all the trouble he might cause there.
+Figure 1 contains an example of a news reel made in our newsroom study with a dialogue between a School District character and a Teachers Union character acting out the news of a teacher shortage.
 As is typical of news reels, the creator is playing both roles.
-Even within the role-play format, framing a news story for social media is challenging.
-In interviews conducted by the Reuters Institute in 2022 with news reel creators (Newman, 2022), many different approaches to creating reels have been found, but a key challenge is to balance between information and entertainment (Newman, 2022).
-If a serious topic is portrayed too light-heartedly, it may diminish the perceived seriousness of the issue (Davis et al., 2022).
-However, if a video does not grab a reader’s attention with something that is interesting and entertaining, they may scroll past it.
-Some news organizations take a conservative approach and stay closer to the TV journalism format, but others lean into the tropes of social media videos to blend news information into an engaging social media style.
-The challenge we address is to help journalists translate their articles into this style, skillfully blending information and entertainment to strike an appropriate tone.
 2.4. ReelFramer AI Tool
 ReelFramer (Wang et al., 2024a) is a creativity support system that uses generative text and image AI to scaffold the process of creating a script and storyboard for a news reel.
-Generative AI is helpful for exploring multiple narrative framings and their details, and generating scripts, character boards, and storyboards.
-The system is designed to be highly interactive, allowing people to make choices and guide the system at every step to ensure the news reels it produces are accurate, authentic, and achievable to execute.
 The user starts by inputting the news article (see Figure 2).
 The system uses a large language model (LLM), GPT-4, to extract locations, people, and activities from the article.
-The user then selects one of the three narrative framings, expository dialogue, reenactment, or comedic analogy.
-Given a framing, the system uses the LLM to suggest foundational details like characters, plot, setting, and key information, which is also called premise in scriptwriting (Batty et al., 2017).
-The user can accept, regenerate, or edit these suggestions.
-Once the user finalizes a premise, the system automatically generates a script.
-Again, the user can accept, regenerate, or edit the script.
-Users are encouraged to explore many narrative framings, premises, and scripts.
+The user then selects one of the three narrative framings, which contain a different balance of information and entertainment (expository dialogue, reenactment, or comedic analogy).
+The system then uses the LLM to suggest foundational details like characters, plot, setting, and key information, which is called the premise in scriptwriting (Batty et al., 2017).
+The system is iterative and built for iteration; users can accept, regenerate, or edit any of the AI generations.
 Once the user is happy with the script, the system provides visual framing to explore the characters’ visual design.
@@
 Based on the character board, it generates a storyboard that shows suggested emotions, actions, and dialogues for each character.
-Users can then make reels based on these materials.
-3. Methods
-3.1. Overview
-We conducted three studies: 1) a 14-week study of how a small newsroom adopts a creativity support tool, 2) an audience reception study assessing the videos made with the tool, and 3) a 4-week study of how two professional TV journalists adopt a creativity support tool.
-We used iterative, exploratory methods (Creswell and Clark, 2017), guided by early insights from student journalists.
-Each study strand was analyzed independently yet informed by prior findings.
-Institutional review boards deemed the project exempt.
-3.2. Organizational Cohort Field Study with Student Journalists
-In the first study, researchers recruited advanced students from a university journalism program that runs a TV and print journalism media outlet called AnonymizedNews to be part of researchers’ simulated newsroom.
-As part of their existing program, the students report local area news, often publishing multiple stories a day with faculty guidance and approval.
-AnonymizedNews  has a fledgling TikTok channel, with fewer than 50 posts and no posts getting more than 3000 views.
+Users can then film the reels based on these materials.
+3. Student Newsroom Study
+To understand the role of human creativity in the presence of an AI support tool, we present a 14-week study of how a student newsroom produces news reels using the ReelFramer AI tool.
+3.1. Methods
+In the student newsroom study, researchers recruited advanced students from a university journalism program that runs a TV and print journalism media outlet called AnonymizedNews.
+As part of their existing program, the students report local area news in print and on television with faculty guidance and approval.
+Before the study, AnonymizedNews had a small TikTok channel, with fewer than 50 posts and no posts getting more than 3000 views.
 Simulated newsroom environments are common in journalism education spaces (Steel et al., 2007) as part of practical and experiential learning with both procedural (environment) and factual (task) authenticity (Chen et al., 2001).
-Choosing students in a university newsroom was fairly representative of professional newsrooms because they often hire recent graduates to run social media, due to their familiarity with social media and the demographic is aspired to reach.
+Choosing students in a university newsroom was fairly representative of professional newsrooms because they often hire recent graduates to run social media, due to their familiarity with social media, and the demographic they aspire to reach (Newman et al., 2022).
 Recruitment emails were sent to senior journalism classes offering an opportunity to learn about creating news reels with AI in a newsroom environment and to regularly publish videos under editorial guidance.
@@
 In this simulated newsroom environment, we had one researcher acting as editor, two researchers acting as market researchers, and five students given news stories each week to convert to reels, as if they were part of a digital news desk.
-The news stories were ones the newsroom had previously reported, sometimes by the student themselves.
-In weekly meetings, researchers acted as both participant observers  (Emerson et al., 2011), taking notes on how student journalists used the tool, and observant participants, interacting with the student journalists to offer the support an editor would provide (Seim, 2024), while allowing space for participants to share insights on production techniques with fellow student journalists as they would in a standard operating newsroom.
-Student journalists were regularly asked about their use or nonuse of the tool in the creation of their newsreels.
-The editor led group feedback discussions for every video, and gave specific feedback for iteration.
+The news stories were ones the newsroom had previously reported.
+In weekly meetings, researchers observed students  (Emerson et al., 2011), and the editor interacted with the student journalists to offer the support an editor would provide (Seim, 2024) in a standard newsroom.
+Student journalists were regularly asked about their use or nonuse of the tool.
+The editor led group feedback discussions for every video and gave specific feedback for iteration.
 After approval by the editor, newsreels were published to the AnonymizedNews TikTok channel.
-3.3. Audience Reception Study
-To compare the quality of the student journalist-produced news reels to the other short form video news content (Shearer, 2024), we conducted qualitative interviews with typical short-form video audience members (Brennen, 2021).
-This allowed us to do two things: 1) access the relative quality of the student production, and 2) provide the student journalists with specific information to improve the audience reception of their videos.
-Questions covered how informative and entertaining the video was, their likelihood to watch the video if they came across it on their personal feed, and the likelihood of the video fitting in with content they typically see, from both qualitative and quantitative perspectives.
-Interviews took between 20 and 66 minutes, with an average of 35 minutes.
-46 individuals signed up to participate, n=42 participants were interviewed.
-Participants were shown six total short-form news videos.
-All six videos were systematically shown in a predetermined balanced design to evenly distribute order effects.
-Participants watched four videos created by the student journalists using ReelFramer on a range of topics.
-Two additional videos were added for comparison: one was professionally created by the Washington Post’s TikTok team to cover the trajectory of a hurricane, the other was created by a student in the AnonymizedNews  program but not in the study and not using the tool; it discussed a local basketball tournament.
-Participants agreed to be interviewed in exchange for course credit as part of the university’s subject pool system.
-All participants were university students from across disciplines aged between 18-30, ($\mu$ = 20).
-Researchers recorded the interviews over Zoom and transcribed the videos before analysis began for axial themes (Saldana, 2003).
-The distribution of quantitative responses was plotted to compare the informativeness and fit of the news reels.
-3.4. Qualitative Cohort Field Study with Practicing Journalists
-To understand if there were differences in how professional journalists would adapt to the creativity support tool, researchers conducted a qualitative cohort field study with practicing journalists.
-In the summer of 2024, researchers recruited a cohort of journalists by posting in a social media group for multimedia journalists.
-Researchers informed participants we would hold four, one-hour collaborative sessions.
-12 individuals completed a demographic questionnaire.
-Four individuals attended at least one training session, two produced at least one video using the tool.
-This sample size and these levels of attrition are not unexpected in longitudinal studies and do not diminish the benefit of the findings (Saldana, 2003).
-Similar to the student study, researchers acted in the capacity of newsroom editor offering critique and advice.
-4. Study 1 Findings: Newsroom Field Study
-4.1. AI Usage: AI as a Creative Springboard
+3.2. Findings: Tool Use and Disuse
 Overall, the student group made 26 news reels, 16 of which were published by the editor.
-The 10 that were unpublished were still being iterated at the end of the session or had been abandoned (Section 4.2.2).
+The 10 unpublished were still being iterated at the end of the session or had been abandoned.
 Altogether, the 16 videos published by the team garnered over 500,000 views.
@@
 13 of 16 published videos used ReelFramer in a significant manner to help achieve an output the editor deemed publishable.
-4.1.1. Reasons for AI Tool Usage
-Creators choose to use the AI tool a majority of the time—for 20 of 26 reels created. Creators who chose to use the tool often described it as “a creative springboard.” There were two main uses.
+Creators choose to use the AI tool a majority of the time—for 20 of 26 reels created.
+To understand what, if any, benefits the tool had, we analyzed why users used the tool and why they chose not to use the tool.
+3.2.1. Reasons for Usage: Framing and Drafting Scripts
+The users repeatedly mentioned the benefit of two of the steps, each supported a different cognitive process. Users mentioned not using the storyboarding features, as it was not cognitively demanding.
 1.
-Selecting the narrative framing and premise.
-They all used the reframing features of ReelFramer to find the characters, premise, and key news facts in the web article that they could use to transform into a role-play script.
+Narrative framing.
+They all used the narrative framing features of ReelFramer to find the characters, premise, and key news facts in the web article that they could use to transform into a role-play script.
 They described this as a cognitively demanding task to “see the story” from a different narrative frame.
-ReelFramer’s suggestions were typically accurate, creating a reliable premise for a script.
+ReelFramers’ suggestions were typically accurate, creating a reliable premise for a script.
 2.
-Drafting multiple scripts.
-They used the scriptwriting writing features of ReelFramer to initiate a draft of the script.
+Drafting scripts from premises.
+They used the scriptwriting features of ReelFramer to initiate a draft of the script.
+Although generating the premise was helpful, there was another cognitive challenge of expanding that to a script.
+Translating raw ideas into fluent text is a complex writing task that requires perspective-taking, thematic grouping and sequencing of information, and managing cohesion and flow  citeflower1981cognitive.
 However, the users consistently reported editing “about half” of the AI-generated script.
@@
 Overall, it was a helpful starting point.
-On the other hand, none of the students used the storyboarding features, as they did not find this part of the task cognitively demanding.
-They said they could “see” how the story would be filmed and apparently did not use the prop, costume, or gesture suggestions—often because they were constrained by their personal set of potential props, costumes, and acting ability.
-4.1.2. Reasons and Risks of Not Using the Tool
-Creators chose not to use the AI tool for 6 of the 26 reels made during the study.
-There were several legitimate reasons not to use the tool.
-However, choosing not to use the tool sometimes resulted in problems.
+Notably, those who used the tool only used it for the narrative framing and scripting steps.
+None of the students used the storyboarding features, as they did not find this part of the task cognitively demanding; they said they could “see” how the story would be filmed.
+3.2.2. Reasons (and Risks) of Disuse
+Creators chose not to use the AI tool for 6 of the 26 reels. In some cases, this was reasonable; in others, it led to problems.
 1.
-Strong understanding of the initial news story.
-Two times, a participant who had used ReelFramer previously chose not to use the tool when making a news reel for a story they had reported.
-Because they conducted the initial research and interviews and framed the original written story, it was easier to understand the key points in the story and imagine it from multiple perspectives, making it cognitively easier to reframe around a role-play narrative.
+Strong understanding of the story (2).
+Two creators skipped ReelFramer when adapting stories they had originally reported.
+Having done the initial reporting made it easier to reframe the story into a role-play narrative without the tool.
 2.
-Task deemed easy.
-P5 made his first news reel with ReelFramer but chose to make the other three without it because he understood the framework well enough to do it on his own.
-However, this decision had trade-offs.
-When the editor reviewed the reels made without ReelFramer, she appreciated several creative additions that the tool would not have suggested (Section 4.2).
-However, the scripts were insufficient for publication—as they all lacked sufficient information about the story.
-The task of creating news reels is often more difficult than it looks; balancing between information and entertainment is hard, and the tool seems to help balance those competing design requirements.
+Task deemed easy (2).
+P5 used ReelFramer once, then stopped, believing he understood the framework.
+While his reels had creative additions and were entertaining, they were not publishable as they all lacked key news information.
+This suggests ReelFramer could help as a scaffolding tool, even for those who do not want or need it.
 3.
-Independent inspiration.
-P2 typically used ReelFramer, but for his fourth video, he had an inspiration for a role-play narrative and developed the framing and script on his own.
-However, he deviated from the role-play framing and wrote the script as a monologue, which the editor and other creators found confusing.
-Once he iterated to return it to role-play form, it was published.
-This incident also shows the risks of not using the scaffolding provided by the tool, although this creator was able to recover.
+Independent inspiration (1).
+P2 had an independent inspiration for his own framing (a monologue rather than a role-play), which he executed by himself.
+The editor liked the inspiration, but found the monologue format confusing.
+He iterated to bring back the role-play style framing and was ultimately successful in publishing the video.
+Independent inspiration could become more scaffolding.
 4.
-AI narratives inappropriate for a serious topic.
-P1 typically used ReelFramer, but for her second video, ReelFramer was not producing a narrative that was serious enough for her topic, which dealt with domestic abuse.
-ReelFramer’s “quirky” and “fun” narratives clashed with the gravity of the topic.
-This clash led her to create the news reel on her own.
-However, she had written the initial news story, so she had a strong understanding of it, and she took inspiration from the tool and followed a role-play format that delivered three news facts.
-The editor approved it for publication, and this news reel received by far the most views.
-There could be many reasons for this success—topic, timing, luck, or unknown quirks of the platform’s sharing algorithm.
-This is discussed in Section 5.
-Despite her success in making a news reel without the tool, she chose to return to using ReelFramer for all her future videos due to the cognitive support it provided.
-Generally, the group was successful in their task, producing 16 news reels garnering over 500,000 views.
-The tool was treated as a creative springboard, assisting with cognitively demanding tasks of reframing a narrative, and writing a script that balanced entertainment and information.
-However, creators did not need a creative springboard when they had independent inspiration or when they had a strong understanding of the original web article.
-Additionally, when the tool failed to provide an appropriate framing or script, they had to fall back on their own creativity.
-Figure 3. Creative additions made by humans that AI does and does not support.
-4.2. Role of Human Creativity: Human Creativity Where AI Left off
+Tone-deaf AI (1).
+P1 skipped the tool when its “fun” and “quirky” tone did not match the seriousness of the topic—domestic violence.
+She took inspiration from the format she had learned from the AI tool and wrote her own script, and the video was successfully published.
+For her future reels, she returned to the tool due to its cognitive support.
+3.2.3. Editors’ Reflections Over Time
+In general, the tool made two cognitively challenging tasks easier—it was not always necessary, but it was helpful.
+The editor observed that in the beginning, the tool served as a teaching instrument to scaffold the process during the learning phase.
+This was particularly helpful for reframing—whereas journalists are taught to write stories in a style that is typically serious, and written from a more objective point of view, the reframing steps that help journalists “see” a story in this new way, also provided the scaffolding to help them find these transformations for themselves.
+Additionally, when a user chose not to use the tool, their reels had fundamental problems such as not balancing information with entertainment—the core challenge the tool supports.
+For some users, positioning the tool only as a scaffolding tool for learning could help them understand the complexities of the task before setting out on their own.
+There were incidents when the tool failed, and it was clear that the creators could make new reels without it.
+However, 4 of the 5 users still chose to use the tool.
+They treated it as a “creative springboard”—rather than a replacement for their creativity, particularly when it came to drafting scripts based on the premise.
+Finding the words to use in a script is hard, and it is easier to edit a script than generate one from scratch.
+3.3. Findings: Role of Human Creativity
+Figure 3. Creative additions made by users.
 To understand when and how creators chose to use their own creativity rather than relying on AI, we analyzed all their creative additions made to news reels created during the study that ReelFramer did not (or could not) make.
-Every week, when the editor reviewed news reels made by the creators (both published and unpublished), she noted all the instances when the creator did something creative outside of what ReelFramer suggested.
-After the study, we analyzed the types of tasks creators choose to innovate on, and found three clear groups of times when creators made creative additions.
-4.2.1. Human Creativity to Improve AI Outputs
-In the study, the editor noted a few tasks that were supported by AI, but for which the creators decided to improve with their own creativity by using their own framing or scriptwriting choices.
-For framing, there were two times when a creator rejected AI character suggestions, and came up with characters that AI would not have: using an inanimate object (a gas pump, to talk about oil prices) or introducing characters not in the story (an Uber Driver, and a passenger to talk about a new bridge).
-For scriptwriting, one creator merged facts from two different web articles about the same topic (a local election) to tell a better story.
-ReelFramer only supported creating scripts from a single web article, so the creator had to add details to the script herself.
-Creators frequently had to edit AI’s attempts at humor in scripts.
-AI often used puns, which creators did not like.
-In the middle of scripts, creators typically removed the attempted humor, but at the end of scripts, creators needed to come up with a new ending like saying something sarcastic.
-One creator added a curse word for emphasis said by a construction worker character.
-AI would never have suggested a curse word, and although the editor made him bleep it out before publishing, the audience still understood the effect, and it was a bold and creative choice.
-Although editing humor was common, the other creative framing and scripting edits were rare—occurring in only 4 of the 26 news reels.
-4.2.2. Human Creativity for Major AI Failures
-Major AI failures were rare—occurring in 3 of the 26 videos, but required significant creativity on the part of creators.
-In all three cases, AI made socially inappropriate suggestions because it lacked context of the story or the creators.
-The first story was about new doula services offered for expecting mothers.
-The original web story featured quotes and reactions from women of color, since they are disproportionately at risk during pregnancies.
-AI suggested roleplaying characters from the story who were pregnant Black women.
-However, the creators did not feel like they could play this in a respectful way.
-But since the service is generally available for all people, they changed the characters to be people indirectly affected by the service (concerned family) and were able to report on the effects for women of all races.
-By creatively changing the characters, the creators were able to reframe the story in a socially appropriate way.
-The second story was about domestic violence, which was a subject that was too serious for the lighthearted tone in the AI-suggested script.
-The creator chose to write the script herself, although she was able to creatively adapt the role-play framing by making the Domestic Women’s Shelter the expert explaining the dangers of emergency alert testing to the “naive” FCC—emergency alerts can expose hidden phones that domestic abuse victims often have.
-The third, and most difficult case was when the team decided it was inappropriate to make a news reel out of a story.
+This analysis is based on a list of creative additions noted by the editor every week during feedback sessions.
+This includes both published and unpublished reels.
+Some changes, like correcting an awkward AI phase, or adding captions, or correcting facts, were not deemed creative, as they were fairly typical copy editing changes.
+However, changes like introducing new types of characters, or adding new elements (actions, personality, animations, etc) were deemed creative because they significantly deviated from the outputs that AI could have produced.
+Across the 26 videos made by the students, the editors noted 22 instances where the creator made substantial creative additions (see Figure 3).
+The editor also denoted the four parts of the creative process where they were made.
+1.
+Framing or Premise (7) Overall, 7 creative additions were made in the framing or premise of the news reel.
+Whereas framing and premise are supported by the AI tool, users choose to make improvements to this area based on either inspiration or need.
+This includes finding different characters to base the role play around, introducing new structures for the narrative, and deciding not to make a news reel at all.
+Whereas the tool limited itself to using characters directly taken from the news story, the users knew that this was not strictly necessary for TikTok videos.
+They could use inanimate objects or tangential characters if they were more fun or easier to depict in the video.
+The tool was also artificially limited in its structures, but users needed to change that—this included merging facts from two news stories to improve the narrative (rather than working with only one story as the tool assumes).
+The tool also assumes the tone should be quirky and fun, but for a story about a serious subject of domestic violence, the creator realized the “fun” tone was inappropriate, and decided to make a role-play but with a serious tone.
+This was a large departure from the tool—and the typical newsreel format—but it was necessary, and the reel got the most views by far.
+Users also needed to exercise creative and professional judgment to determine when a news reel format would be inappropriate.
 The article was titled “Women’s Club Basketball Discusses What Black History Month Means to Them”.
-None of the journalists identified as both Black and female, and thus role-play style narrative felt socially inappropriate.
-Additionally, they did not want to reframe the narrative around a different set of characters because it would displace the voices of the very people it was trying to center.
+None of the journalists identified as both Black and female, and thus role-play narrative felt socially inappropriate.
+Additionally, they did not want to reframe the narrative around a different set of characters because it would displace the voices of the people it was trying to center.
 Ultimately, they decided that not making the reel was the right decision.
-Although major AI failures were relatively rare, they required human creativity to detect and address them.
-4.2.3. Human Creativity Where AI Leaves off
-In the study, the editor noted many tasks that were not supported by AI that creators decided to add their creativity to.
-This happened in 12 of 26 news reels (Figure 3).
-ReelFramer was designed to address the full creative process from framing to storyboarding.
-However, the creative process has no bounds, and the creators found many tasks to do creatively that the tool did not address.
-There were four instances of creators adding personality to the characters, which was largely shown through acting ability.
-For example, making a teenager have an aggressively dismissive attitude about a new college program, to bring drama and realism to the character (as well as making it sound less like an advertisement).
-There were three instances of using props in highly effective ways.
-First, using a real dog or baby instead of a doll, and second, in an episode about a soup kitchen, had a chef-like character actually making a dish and using a different cooking utensil in every scene to show progress.
-None of these were suggested by AI, but were things creators realized they had on hand.
-The market research done in week 10 (Section 5) revealed that the videos would be more interesting if there were more movement, especially at the beginning.
-Thus, in the last week, creators added two more actions to videos – pouring maple syrup, jogging, and unwrapping food.
-There were four other notable creative additions, like using animations of a truck for a story on monster trucks, adding sound effects, and adding colored lights and captions to reflect elements in a story.
-Although the AI tool attempted to cover the entire creative process, creators found new dimensions where the tool left off, and they were able to add many of their own creative touches.
-4.3. AI Risks: Need for Training Humans to Critique due to AI Overreliance
-AI is widely known to produce errors—as are humans.
+2.
+Script (2) Overall, most users appreciated AI most for its ability to produce scripts.
+All AI-generated scripts needed editing, but sometimes those edits had to be creative.
+AI would often add terrible jokes to scripts and users had to find a way to improve the humor—sometimes they deliver the AI joke sarcastically, to make it funny (which is not how it was intended).
+Sometimes, lines were dull and needed more energy.
+One user added a curse word, which the editor insisted he bleep, but it still brought more energy and was funny.
+In both cases, users brought their own sense of humor to the script.
+3.
+Reel Production (10) A major area where users added creativity was in the production of the reel—acting, props, lighting, sound effects, and post-production like actions and animation.
+This is an area that the tool did not support well, and that users found there were many ways to enhance the quality of reels—as well as bring in their own knowledge and creativity.
+Adding a strong personality to the characters was a creative decision that made a big difference.
+AI scripts had a negative tendency to sound too positive or even promotional.
+But creators could take an enthusiastic character and make them sarcastic, or have them be rudely cut off by another character to undermine the overzealous positivity.
+Giving characters amplified emotions (such as being very surprised, or very miserable when sick) also gave videos more life.
+Using props well also made a big difference, and creators used what they had on hand—including a real dog for a story about animal shelters, or cooking utensils for a story about food insecurity.
+AI could not have known what users would have, so its suggestions for props were weak, but users reacted to their environment to add drama through prop usage.
+Additionally, CapCut software makes post-production effects like animations, sound effects, lighting filters, and caption colors very easy.
+Users often found creative ways to exchange the message through this type of post-production.
+4.
+Iterating from Audience Feedback (3) In the last week, creators added three more actions to videos—pouring maple syrup, jogging, and unwrapping food.
+Each of these was highly relevant to the story and made the videos more fun and eye-catching.
+Overall, users made significant creative additions to their reels.
+This included adding to processes the tool supported, such as framing and scripting, as well as places where AI left off—such as acting, visual effects, and props, or iterating based on user feedback.
+When making these additions, users often injected their own knowledge and taste, like their own sense of humor, their acting ability, their knowledge of post-production techniques, and the props they had on hand.
+Moreover, this level of creativity indicates that users were not passively accepting AI’s creativity.
+They actively inserted themselves in the process based on need (when AI failed) or inspiration (when they had independent ideas—often that were better than AI’s suggestions).
+3.3.1. Editors’ Reflections Over Time
+To the editor, it was clear that the creators were heavily invested in producing high-quality videos.
+Whereas AI could produce many outputs, the creators took ownership over the entire process.
+They critiqued outputs, considered the audience, adjusted based on what was appropriate and available, and generally tried to maximize the creative potential of each story, within the time constraints of their deadline.
+In stark contrast, AI was mostly following a template and executing a task as directed.
+The creators’ sense of ownership and holistic responsibility for the quality of the product drove their continued creativity.
+3.4. Overcoming AI Errors
+Major Issues (15)
+AI Attributed (11) Human Attributed (4)
+too positive/promotional (4)
+too long (2)
+too few facts (2)
+unclear narrative (1)
+too wordy (1)
+abrupt intro (1)
+slow paced (2)
+confusing format – monologue (1)
+biased narrative (1)
+Table 1. Major Issues: AI vs Human Attributed in the student newsroom study.
+Minor Issues (19)
+AI Attributed (15) Human Attributed (4)
+awkward AI phrase (4)
+confusing character names (3)
+ending lacks polish (2)
+too positive/promotional (1)
+strays from news (1)
+tone shift (1)
+unclear/missing fact (2)
+unnecessary dialogue (1)
+labels too long (1)
+added curse word (1)
+add relevancy to the present (1)
+missing facts (1)
+Table 2. Minor Issues: AI vs Human Attributed in the student newsroom study.
+AI is widely known to produce errors—as do humans.
 The newsroom in this study had an editor who was in charge of quality control.
 This is typical of all newsrooms.
-Similar to typical editorial newsroom dynamics, creators submit drafts, receive critique and edits from the editor and have to iterate to produce a new draft.
-Iteration continues until the editor decides to publish the piece, or decide not to run it.
-To understand the nature of the edits, we tracked two things: the magnitude of the error (major, minor or none), and the attribution of the error (whether the error was introduced by the AI or the human journalist).
-For error magnitude, major edits are those that required reshooting the entire news reel; minor edits required reshooting only part of the video (e.g., a localized scene or dialog edits); no edits meant it was ready for publication.
-Table 3 shows the magnitude of edits for each news reel’s first iteration.
+Similar to typical editorial newsroom dynamics, creators submit drafts, receive critiques and edits from the editor, and have to iterate to produce a new draft.
+Iteration continues until the editor decides to publish the piece or decides not to run it.
+To understand the nature of the edits, we tracked two things: the magnitude of the error (major or minor) and the attribution of the error (whether the error was introduced by the AI or the human journalist).
+For error magnitude, major edits are those that required reshooting the entire news reel; minor edits required reshooting only part of the video (e.g., a localized scene or dialog edits).
 We also tracked whether the errors were introduced by AI or by the creators, as judged by the editor from discussions with the creator.
@@
 These errors were accepted by the creators, and they filmed news reels with these errors in them.
-Although it is expected that AI will make errors, and that many people using AI on a deadline might simply accept AI content at face value, skimming the content and not checking it carefully.
-However, this case is different—creators actually had to speak the lines AI wrote, so they were fully cognizant of the dialog they were including.
-This was true for both student journalists and professional journalists (Section 6).
-It was surprising to the editor because being critical is an essential aspect of journalism training.
-Major Issues (15) AI Attributed (11) Non-AI Attributed (4) too positive/promotional (4) slow paced (2) too long (2) confusing format - monologue (1) too few facts (2) biased narrative (1) unclear narrative (1) too wordy (1) abrupt intro (1)
-Table 1. Major Issues: AI vs Non-AI Attributed issues in news reels created in the newsroom study.
-Minor Issues (19) AI Attributed (15) Non-AI Attributed(4) awkward AI phrase (4) labels too long (1) confusing character names (3) added curse word (1) ending lacks polish (2) add relevancy to the present (1) too positive/promotional (1) missing facts (1) strays from news (1) tone shift (1) unclear/missing fact (2) unnecessary dialog (1)
-Table 2. Minor Issues: AI vs Non-AI Attributed issues in news reels created in the newsroom study.
+Although it is expected that AI will make errors, and that many people using AI on a deadline might simply accept AI content at face value and not check it carefully.
+However, this case is different—creators actually had to speak the lines AI wrote, so they were fully cognizant of the dialogue they were including.
 The most common major issue was clearly attributed to the AI tool—sounding too positive.
 The AI is trying to balance entertainment and information in a “quirky” way, and the stories should not sound sad and serious like traditional evening news anchors.
-However, the AI leaned too heavily toward a positive tone and this bias clashed with the content in different ways.
+However, AI leaned too heavily toward a positive tone, and this bias clashed with the content in different ways.
 For stories that were serious or sad (stories about homelessness, domestic violence, or climate change), the “quirky” tone was inappropriate for the topic.
-For many lighter stories—like a small town winning an award for being charming—AI’s positive tone was also a problem.
+For lighter stories—like a small town winning an award for being charming, AI’s positive tone was also a problem.
 The positive tone made the video feel like an advertisement rather than news.
@@
 These stick out badly and are fairly localized changes to make, so the effort to change them should have been minimal.
-Phrases like “a little bird told me’’ are awkward AI phrases that were inappropriate for a news reel because all news should make some effort to attribute information to sources.
+Phrases like “a little bird told me” are awkward AI phrases that were inappropriate for a news reel because all news should make some effort to attribute information to sources.
 Other awkward phrases include “a comedy of wonders” and distracting attempts at wordplay such as “a two-tiered system?
-is that kind getting rewarded with two kinds of deserts?” When asked why they included these phrases, the typical response was “I guess it seemed OK.”
-This acceptance indicates an implicit trust in the AI, even when AI does a task in the creator’s realm of expertise.
+Is that kind getting rewarded with two kinds of desserts?” When asked why they included these phrases, the typical response was “I guess it seemed OK.”
+This acceptance indicates an implicit trust in AI, even when the AI does a task in the creator’s realm of expertise.
 A crucial role of human creativity is to be critical of AI outputs.
 Even journalists who are trained to be critical may require additional training to learn to be critical in this new way.
-4.4. Human Creative Role Over Time
-We also tracked the magnitude and attribution of errors in the first iteration of all 26 news reels to track errors over time.
-As expected, every creator’s first video required major edits, but later videos needed more minor edits, or even no edits.
-Participant Article 1 Article 2 Article 3 Article 4 Article 5 Article 6 P1 Major Minor Minor Minor No edits No edits AI errors 2 1 1 1 Non-AI errors 1 1 P2 Major Minor Major Major Major Minor AI errors 3 1 3 1 Non-AI errors 1 1 P3 Major Minor Minor Major / / AI errors 3 2 2 2 / / Non-AI errors / / P4 Major Minor Major Minor Minor No edits AI errors 1 1 2 1 1 Non-AI errors P5 Major Major Major Minor / / AI errors 2 / / Non-AI errors 1 1 1 1 / /
-Table 3.
-Classification of the amount of editing needed for news reels made for each article and each participant.
-Edits were either major (requiring a full reshoot), minor (localized scene or line edits) or no edits.
-Errors are broken into those introduced by AI (and accepted by creators without realizing), or errors not introduced by AI.
-Cells filled with a slash indicate creators did not make news reels (some participants made more than others).
-4.4.1. Learning the Limits of the Tool, and Reasserting Professional Standards
-Early news reels were problematic because the news conveyed in them was unclear.
-This was likely due to a combination of problems: being too long, having meaningless dialog lines, and not containing enough facts.
-Additionally, when the tone was too positive, the reels felt like ads, not news, and when the roles of the characters were unclear, the video was hard to follow.
-All these problems could be attributed to AI, and creators’ early over-reliance on the tool.
-Through conversations with the editor and each other, they were able to overcome this.
+3.4.1. Editors’ Reflection Over Time
+The editor noticed a general trend that, initially, users relied heavily on the AI’s outputs—even accepting its blatant errors without question.
+However, through critical reflection and iteration, creators were able to reassert themselves.
+AI was initially seen as an authority on news reels, due to the creators’ lack of experience and AI’s mystique.
+But ultimately, users were better able to balance AI’s authority and their own.
+Through conversations with the editor and each other, creators were able to overcome this.
 It seemed to be correlated with both more group and editorial critique, as well as students having a better mental model of the strengths and weaknesses of the tool through repeated usage.
 The editor pointed out that “at some point [the students] were able to anticipate what the tool was going to suggest (to a point) because they understood how it was translating between genres [web stories to reels].”
-In the early videos, students often “miss the heart of the story.” For example, in a video about falling gas prices, the student “missed the main connection between the winter months and the change in gas prices”.
-Many exchanges in the early videos are wordy and contentless and do not help convey the news.
-It was important for students to understand the “news value” so the videos focused on the information rather than entertainment.
-The editor said that the students’ “learning/improvement happened when they made connections between the tool and their occupational standards.” Balancing between news and entertainment was one of the goals of the tool, and whereas it did help, it was still important for the creators to have their own understanding of this, so they could make good edits to what the tool was suggesting.
-4.4.2. Learning from market research
-In week 10 of the newsroom study, the team was given a report from the market research team detailing ways their news reels could be improved.
-This included more attention-grabbing elements early in the video, and more character motion-like actions within the scene (Section 5).
-These suggestions provided valuable insights because these were all emerging trends appearing in non-news-based social media videos that had not yet become integrated with news reels and were imbued into ReelFramer.
-Since news reels have to compete for attention with all the other reels, they will eventually have to adopt these trends to maintain their audience.
-Working as a team, the creators brainstormed relevant actions they could put in their current videos and found several compelling ways to do this.
-For an article about tapping maple syrup trees early due to global warming, the video started with the character pouring syrup on waffles while talking.
-Another video about the opening of baseball season had several characters using props and doing simple, yet interesting actions like unwrapping food.
-A third added fun animation of a truck for a story about a monster truck rally.
-The market research feedback was insightful, relevant and easy to apply. It was also outside the scope of AI, which cannot keep up with emerging trends.
-5. Study 2 Findings: Audience Reception
-Figure 4.
-Violin plots showing the distribution of ratings of videos as entertaining on a 1-7 scale.
-P is the professionally-created video, S the student-created video without using the tool and RF1-4, four videos created using the tool.
-The small letters labelling the bars identify groups of videos that are not statistically significantly different according to Tukey’s HSD.
-Multiple letters indicate that video is in two groups.
-Specifically, there is not a significant difference in the rating between videos P, S, RF2 and RF4, nor between video S and RF1-4 but the differences between video P and RF1 and RF3 are significant.
-Figure 5. Violin plots showing the distribution of ratings of videos as entertaining. Labels are the same as Figure 4.
-The main challenge in a news reel is to balance information and entertainment.
-We test how informative and entertaining news reels created in the study (using ReelFramer) were, compared to a popular professional news reel, and a news reel produced by a student at the same university but who was not in the study and was not exposed to ReelFramer.
-First, we assessed how entertaining the news reels made during the story with ReelFramer were compared to the two baseline reels.
-According to a one-way ANOVA test, there was a significant difference in how entertaining the videos were ($F(5,229)=3.3$, $p=0.007$).
-Half of the videos produced using the tool were rated as as entertaining as the professionally made video.
-The non-study student’s video was also as entertaining as the popular video (Figure 4).
-However, it was rated less information than 3 of 4 ReelFramer videos.
-Next, we assess how informative the news reels produced in the study are to the popular video and the non-study student reel.
-A one-way ANOVA test (not assuming equal variance in the groups) shows participant ratings of the video’s informativeness differed significantly ($F(5,108.18)=7.7$, $p<0.001$).
-Three-quarters of news reels produced in the study using ReelFramer were rated as “as informative as” the professionally produced video (Figure 5).
-These results indicate that ReelFramer reels successfully balanced information and entertainment much of the time.
-They were competitive with the popular professional news reel, and it was better than the non-study student reel because they were more informational.
-To assess audience acceptance of the news reels, we asked subjects whether they would watch each of the videos.
-Roughly three-quarters of subjects said they would watch the professionally produced video, 56% would watch the videos created with the tool and 45% would watch the student video produced without the tool.
-This difference was statistically significant ($\chi^{2}(5)=14.6$, $p=0.01$).
-Roughly half said they would share the professionally produced video, 40% would share the study videos, and only a quarter would the non-study student video.
-This difference was statistically significant ($\chi^{2}(5)=11.6$, $p=0.04$) There was no statistically significant difference in the perceived “fit” of videos on participant feeds ($\chi^{2}(5)=2.52$, $p=0.77$), though the student-produced video had a lower fit than other videos.
-Qualitative data suggests that this has to do with the subjects they prefer to watch (pets, sports, etc.) rather than quality.
-These results suggest that the videos produced with ReelFramer were closer to the professionally produced video in terms of acceptability than the non-study student video.
-After conducting audience reception interviews, researchers identified strategic and manageable production changes they suggested to student journalists.
-Audience members indicated they would like more attention-grabbing elements early in the video and adding moving character movement within the scene.
-Student journalists wove this feedback into their subsequent productions and asked for more from the tool to match audience expectations.
-Similar to the way a newsroom would operate, journalists had to option of bringing market research into their production, continuing to iterate on their creative process.
-6. Study 3 Findings: Professional Journalists
-To supplement the results of the newsroom study done with student journalists, we conducted a 4-week study with two professional TV journalists from local news stations in the US.
-6.1. Reasons for AI Tool Usage
-Both journalists made 3 news reels each, for a total of 6 newsreels.
-They both chose to use the AI tool for all their videos.
-Similar to the newsroom study, they used the AI tool for framing and script writing, but not for storyboarding—they described framing and script writing as feeling “very different” than their training, and difficult to “get started with.” They also said that they had to edit “about half” the AI scripts for clarity, correctness, and smooth language.
-6.2. Role of Professional Human Creativity: Human Creativity Where AI Leaves off
-In four of the six videos, the journalists made notable creative additions.
-This included some of the same categories seen in the student study—adding personality to characters through acting and adding visual effects.
-However, they also introduced new elements.
-Two videos included real footage as background imagery.
-One video used an audio track.
-One video took the role-play premise to an extreme and introduced a time-lapse—where a citizen called city services over the course of 10 years to get a fallen tree removed (the city representative promised to “get back to them” in a robotic voice in every call).
-With professional journalism training and resources, it is possible to add even more creativity where the AI tool leaves off.
-Of the six videos in the study, none experienced the major AI failures that forced students to make the news without the tool.
-However, AI did create some problems.
-Participant Pro 1 made a video discussing a new scam.
-The character introducing the scam was the Better Business Bureau (BBB).
-Although the typical TV news audiences know what the BBB is, Gen-Z and even Millennial audiences are less familiar with and are confused about the connection between the BBB and scams.
-AI was not aware of this gap in audience knowledge, but the editor was.
-But that decision still required much discussion and consideration.
-6.3. AI Risks: Even Professionals Need to Hone Critique of AI
-Major Issues (6) AI Attributed (4) Non-AI Attributed(2) unnecessary dialog (2) slow paced (1) too few facts (1) added timeline narrative unclear (1) too long (1)
-Table 4. Major Issues: AI vs Non-AI Attributed issues in news reels created in the professional study.
-Minor Issues (8) AI Attributed (5) Non-AI Attributed(3) confusing character names (3) distracting music (2) awkward AI phrase (2) should not name specific unrelated brands (1)
-Table 5. Minor Issues: AI vs Non-AI Attributed issues in news reels created in the professional study.
-Despite years of journalistic practice, the professionals still made many errors.
-Four of the six major errors and five of eight minor errors were introduced by AI and not caught by professionals before filming (see Tables 4 and 5).
-Generally, they were the same types of errors seen in the student study.
-This included major errors like having unnecessary and contentless dialog (“I can’t thank you enough”, “We’re so happy to help”), having too few facts, and being too long.
-As well as minor issues, like using awkward AI phrases (“How about I come and start training my own railings?”).
-The editor thought these should have been glaring errors to catch.
-Upon viewing the iteration of the video containing that dialog, the editor asked if “training railings” was a concept the professional journalist knew (perhaps a local phrase).
-The journalist said no, “The AI said it; I thought it was a thing.” Despite years of journalism training—a field that prides itself on being critical—these professionals were imperfect AI critics.
-6.4. Professional Creative Role Over Time
-Similar to the student study, news reels made by professional journalists still required major edits.
-However, both professionals required no edits by their third video (see Table 6).
-In both videos needing major edits, the errors were similar to the students’—the news content was unclear due to a variety of errors, like contentless dialog, and unclear roles.
-They did have to be coached by the editor to correct AI, but once the editor pointed that out, they adapted almost immediately.
-This is likely because they had a wealth of journalism training and understanding of “news value”, and thus faster to integrate their news training into their use of the tool.
-Participant Article 1 Article 2 Article 3 Pro1 Major Minor No edits AI errors 1 3 Non-AI errors 3 1 Pro2 Minor Major No edits AI errors 1 3 Non-AI errors 1
-Table 6. Classification of the amount of editing needed for news reels made by professional journalists.
-7. Discussion and Conclusion
-7.1. Experts Value AI Creative Support and Use Their Creativity Where AI Leaves Off
-We found that when AI presents a solution, that solution, unless obviously flawed, is generally adopted.
-The user then applies their creativity to the parts of the project not solved by AI.
-There may be something to be learned from other areas of design.
-Lars Lerup, in a book called Building the Unfinished, observed that the physical environment will change over time: for example, we renovate buildings (Lerup, 1977).
-Instead of designing something that is a finished product, it might be better to build it in a way that encourages or facilitates later modification.
-A contemporaneous and related idea from architecture has to do with planning and building incrementally rather than all at once (Alexander, 1977).
-AI leaves many things unfinished.
-In the newsroom study, AI failed in spectacular ways, leaving people to complete the task.
-AI also does not know what the next trend will be.
-In the newsroom study, we saw how the results from the market research were immediately used by journalists to keep up with social media engagement trends that AI was unaware of.
-Creativity is about discovering what comes next, and so AI has not shown abilities to do that.
-With all automated tools there is a fear that human skills will atrophy from disuse.
-However, in this study, we saw rampant use of creativity, especially where the tool left off.
-This incompleteness of AI could be thought of as a feature rather than a bug—it is an opportunity to keep us engaged with the tool.
-Future work might explore the different ways of leaving an artifact unfinished and their effects on the quality of the end product, the overall level of productivity, and the long-term effect on human skill.
-7.2. Creating Vigilance to AI Errors
-AI makes errors.
-Often it makes obvious errors like using totally made up phrases.
-It was surprising to see that journalists, even professionals, filmed videos without fixing them.
-There seems to be a pervasive assumption that AI is an expert, even when AI is operating in a field in which the human is the expert and the AI is meant to play a supporting role.
-Whatever the reason for their deference, people need to develop a vigilance towards AI errors.
-AI is so easy to generate that it seems to lull us into System I reasoning (Kahneman, 2011), where we think fast and mostly go with the flow.
-However, when working with AI, we need to develop ways to help people shift into System II reasoning, where we think slowly and think critically.
-System I is our natural state because it will take more than just willpower to develop this vigilance.
-Teams and organizations will have to design ways to avoid falling for AI errors.
-Editors are needed to be critical; editors will need to learn the new flaws that AI is introducing into creative work.
-For example, AI is known to lean toward cliches in creative domains  (Chakrabarty et al., 2024; Mirowski et al., 2024), known to make up phrases (as shown here), known to lack social context, and known to be bad at simple math (Schick et al., 2023).
-When possible, it might also be helpful to constantly track what text or images were AI-generated so editors know where to focus their attention.
+In the early videos, students often “miss the heart of the story.” It was important for students to understand the “news value,” so the videos focused on the information rather than entertainment.
+The editor said that the students’ “learning and improvement happened when they made connections between the tool and their occupational standards.” Balancing between news and entertainment was one of the goals of the tool, and whereas it did help, it was still important for the creators to have their own understanding of this, so they could make appropriate edits.
 3.5. Limitations and Future Work
-These studies attempt to explore AI at work in a newsroom, but we only studied one news team—and it was one that we assembled for this purpose.
-Although it had many of the traits of a real small-town newsroom, there are limitations in how much we can generalize to all newsrooms.
-Although the study lasted 14 weeks, there is no telling what would have happened if the study had continued after those 14 weeks.
-While we did not see any skill atrophy in the 14-week period, perhaps we would have after a year.
-This paper focused on one particular tool for one particular social media phenomenon, but different tools and trends may have different benefits and challenges.
-However, factors such as accepting AI errors seem to be pervasive problems.
-ReelFramer as a tool is certainly not perfect.
-If its workflow had been better, perhaps creators would not have had so many opportunities to be creative when it failed.
-However, the most important failures were the ones when the new reel form did not fit an article—which was not a property of the tool, but of the narrative form.
-It would be interesting future work to see if people were still using the tool in a year or whether they had embodied the news reel form so well the tool no longer brought them cognitive benefit.
-Or perhaps they might even find it too limiting over time, as they started to understand the design space and perhaps start to innovate on it.
-It would also be interesting to study how the team might improve the tool for themselves.
-Since the tool is basically a structured prompting technique, the creators could easily edit the tool by simply editing the prompts in the text.
-Another form of creativity would be to use prompts within the tool to innovate on the form.
+This study examined one student newsroom using a specific AI tool for 14 weeks, limiting generalizability.
+Future work could explore longer-term effects on populations of multiple types of journalistic and AI-based experiences.
+In particular, professional newsrooms may face different dynamics given their tighter deadlines, higher editorial standards, and larger audiences, and other creative fields may reveal distinct patterns of human-AI collaboration.
+4. Conclusion
+This paper studied the benefits and challenges of using AI in the organizational context of a newsroom.
+AI has many strengths as a creativity support tool—especially for writing—so we studied journalists, experts in the craft of writing and storytelling.
+Newsrooms and other creative industries are places where AI could significantly impact work—either by changing its nature or potentially displacing it.
+Overall, we found that AI was helpful both as scaffolding during the learning phase and as a creative springboard once users understood the task.
+AI supported human creators but did not replace them.
+While it could generate outputs, it was the creators who took ownership of the task, using judgment and creativity to tailor stories for audiences, rather than merely adapting them to templates.
+Despite the flexibility of generative AI, it follows instructions linearly and struggles to adapt to the nuances of individual stories or audiences.
+Because AI still makes frequent mistakes and lacks contextual understanding, editorial oversight was essential to ensure quality and avoid blindly accepting flawed outputs.
+Through this reflective process, creators also reasserted their expertise.
+While AI was initially seen as an “authority,” users ultimately learned to balance its contributions with their own.